# The Landscape of Agentic Reinforcement Learning for LLMs: A Survey

### v1

@@ 6.2 Scaling up Agentic Training
A nascent yet critical frontier for Agentic RL involves a paradigmatic shift from treating the training environment as a static entity to viewing it as a dynamic and optimizable system. This perspective addresses a core bottleneck in agent development: the scarcity of interactive, adaptive environments and the difficulty of engineering effective reward signals. As a growing consensus holds that prevalent environments like ALFWorld [456] and ScienceWorld [458] are insufficient for training general-purpose agents [540], research is moving beyond solely adapting the agent’s policy. Instead, a co-evolutionary approach uses learning-based methods to adapt the environment itself. One key strategy is to automate reward function design. This involves deploying an auxiliary "explorer" agent to generate a diverse dataset of interaction trajectories, which are then used to train a reward model via heuristics or preference modeling. This effectively decouples agent training from the expensive process of manual reward specification, enabling the learning of complex behaviors without direct human annotation.
@@
Beyond automating the reward signal, a second, more dynamic strategy is to automate curriculum generation, transforming the environment into an active teacher. This approach establishes a feedback loop where an agent’s performance data, highlighting specific weaknesses, is fed to an “environment generator” LLM. As exemplified by EnvGen [541], this generator then procedurally adapts the environment’s configuration, creating new tasks that specifically target and remedy the agent’s deficiencies. This form of goal-directed Procedural Content Generation (PCG) ensures the agent is consistently challenged within its “zone of proximal development,” accelerating learning and preventing overfitting. Together, automated rewards and adaptive curricula create a symbiotic relationship between the agent and its environment, establishing a scalable "training flywheel" that is essential for the future of self-improving agentic systems.
@@ 7 Conclusion
This survey has charted the emergence of Agentic Reinforcement Learning (Agentic RL), a paradigm that elevates LLMs from passive text generators to autonomous, decision-making agents situated in complex, dynamic worlds. Our journey began by formalizing this conceptual shift, distinguishing the temporally extended and partially observable MDPs (POMDPs) that characterize agentic RL from the single-step decision processes of conventional LLM-RL. From this foundation, we constructed a comprehensive, twofold taxonomy to systematically map the field: one centered on core agentic capabilities (planning, tool use, memory, reasoning, self-improvement, perception, etc.) and the other on their application across a diverse array of task domains. Throughout this analysis, our central thesis has been that RL provides the critical mechanism for transforming these capabilities from static, heuristic modules into adaptive, robust agentic behavior. By consolidating the landscape of open-source environments, benchmarks, and frameworks, we have also provided a practical compendium to ground and accelerate future research in this burgeoning field.
Acknowledgments
We acknowledge Zhouliang Yu and Minghao Liu for their guidance and discussion.

### v4

@@ 6.3 Scaling up Agentic Environments
A nascent yet critical frontier for Agentic RL involves a paradigmatic shift from treating the training environment as a static entity to viewing it as a dynamic and optimizable system. This perspective addresses a core bottleneck in agent development: the scarcity of interactive, adaptive environments and the difficulty of engineering effective reward signals. As a growing consensus holds that prevalent environments like ALFWorld (Shridhar et al., 2021) and ScienceWorld (Wang et al., 2022) are insufficient for training general-purpose agents (Zheng et al., 2025f), research is moving beyond solely adapting the agent’s policy. Instead, a co-evolutionary approach uses learning-based methods to adapt the environment itself. One key strategy is to automate reward function design. This involves deploying an auxiliary "explorer" agent to generate a diverse dataset of interaction trajectories, which are then used to train a reward model via heuristics or preference modeling. This effectively decouples agent training from the expensive process of manual reward specification, enabling the learning of complex behaviors without direct human annotation.
@@
Beyond automating the reward signal, a second, more dynamic strategy is to automate curriculum generation, transforming the environment into an active teacher. This approach establishes a feedback loop where an agent’s performance data, highlighting specific weaknesses, is fed to an “environment generator” LLM. As exemplified by EnvGen (Zala et al., 2024), this generator then procedurally adapts the environment’s configuration, creating new tasks that specifically target and remedy the agent’s deficiencies. This form of goal-directed Procedural Content Generation (PCG) ensures the agent is consistently challenged within its “zone of proximal development,” accelerating learning and preventing overfitting. Together, automated rewards and adaptive curricula create a symbiotic relationship between the agent and its environment, establishing a scalable "training flywheel" that is essential for the future of self-improving agentic systems.
6.4 The Mechanistic Debate on RL in LLMs
Two competing explanations have emerged for why RL appears to boost LLM reasoning. The “amplifier” view holds that RL with verifiable rewards—often instantiated via PPO-style variants such as GRPO—mainly reshapes the base model’s output distribution: by sampling multiple trajectories and rewarding the verifiably correct ones, RL concentrates probability mass on already-reachable reasoning paths, improving pass@1 while leaving the support of solutions largely unchanged; consistent with this, large-k pass@k analyses often find that the base model eventually matches or surpasses its RL-tuned counterpart, suggesting elicitation rather than creation of capabilities, and further evidence indicates that reflective behaviors can already emerge during pre-training (Shao et al., 2024b; Yue et al., 2025a; AI et al., 2025). By contrast, the “new-knowledge” view argues that RL after next-token prediction can install qualitatively new computation by leveraging sparse outcome-level signals and encouraging longer test-time computation: theory shows that RL enables generalization on problems (e.g., parity) where next-token training alone is statistically or computationally prohibitive; empirically, RL can improve generalization to out-of-distribution rule- and visual- variants, induce cognitive behaviors (verification, backtracking, subgoal setting) that were absent in the base model yet predict self-improvement, and in under-exposed domains even expand the base model’s pass@k frontier (Guo et al., 2025a; Tsilivis et al., 2025; Chu et al., 2025b; Gandhi et al., 2025; Cheng et al., 2025). Whether RL can truly endow LLMs with abilities beyond those acquired during pre-training remains an open question, and its underlying learning mechanisms are still to be fully understood.
Case study: Mathematical Reasoning
From a mechanistic standpoint, our survey of RL for mathematical reasoning in Sec 4.3 suggests that RL functions neither as a pure “sampler amplifier” nor as a universally reliable source of genuinely new reasoning algorithms (Yue et al., 2025a). Across the cited mathematical and code-reasoning studies, approximately 2/3 primarily emphasize improvements in pass@1 accuracy, while about 1/3 explicitly report expanding pass@k frontiers (e.g., higher pass@32 at fixed or only modestly improved pass@1), indicating that many systems leverage RL chiefly to reshape the sampling distribution over pre-existing competent trajectories rather than to unlock qualitatively new ones. However, cases such as 1-shot RLVR and self-evolving System-2-style frameworks (e.g., rStar-Math–like pipelines (Guan et al., 2025a)) also exhibit “post-saturation” generalization and cross-category transfer, which are difficult to explain as mere reweighting and instead suggest strategy-level reorganization of latent capabilities.
Empirically, we find that such “new-capability” behaviors appear most reliably on tasks with (i) high-fidelity, often executable or formally checkable reward signals; (ii) compositional or multi-step structure where many partial trajectories are verifiably graded; and (iii) base models in the “intermediate” regime (neither near-random nor near-ceiling) where the space of near-miss trajectories is rich enough for exploration but still densely populated with correct reasoning paths. Under these conditions, policy-gradient updates plus explicitly managed exploration (e.g., entropy bonuses, self-play curricula, or search-guided expert iteration) seem to move the model toward internalizing more abstract decision rules—whereas on easier, low-noise benchmarks or with coarse outcome-only rewards, RL predominantly acts as an amplifier that sharpens and reuses patterns already implicit in the pretrained model.
6.5 Architectural Patterns for Real-World Agent Deployment
While the survey primarily analyzes RL as a mechanism for improving reasoning performance, the practical deployment of RL-optimized systems requires architectural patterns that ensure reliability, safety, and operational robustness. This subsection synthesizes four cross-cutting design principles—safety guardrails, human-in-the-loop supervision, hierarchical orchestration, and inter-agent communication protocols—that commonly arise in real-world deployments of RL-enhanced reasoning systems, irrespective of the domain.
Guardrails and Safety Patterns.
Deployed systems typically incorporate multi-layered safety mechanisms that operate independently of the RL optimization loop. These include input validation (schema enforcement, semantic filtering, and constraint checking), output sanitization (format normalization, groundedness checks, and post-hoc constraint satisfaction), and sandboxed execution for tool or code calls. Such guardrails can be implemented in two major ways: (1) Using RL optimization itself as a safeguard, where, for example, many works directly incentivize models to “think safely” during the reasoning output via RL (Zheng et al., 2025c; Zhang et al., 2025z) ; and (2) Using external modules to monitor RL training, such as AWS Bedrock.
Human-in-the-Loop Verification.
Human oversight remains essential in high-stakes or uncertainty-prone settings (Mozannar et al., 2025; Takerngsaksiri et al., 2025). HITL mechanisms range from synchronous review of critical decisions to asynchronous auditing, exception handling, and feedback collection. They often rely on model confidence signals or external uncertainty detectors to trigger intervention (Nazir & Banerjee, 2025). Architecturally, HITL provides sparse but high-fidelity corrective signals that complement RL reward structures, enabling safe deployment even when real-world reward feedback is limited, delayed, or noisy.
Hierarchical Orchestration.
Many practical systems adopt hierarchical control structures (such as supervisor–worker, controller–executor, or planner–solver patterns, as observed in (Zhang et al., 2025r; Liu et al., 2025c; Hu et al., 2025c)) to manage complex workflows. The supervisory layer coordinates subtasks, resolves conflicts, or enforces global constraints, while lower-level components focus on domain-specific reasoning or tool execution. This decomposition facilitates temporal and structural credit assignment, improves scalability, and mirrors enterprise orchestration pipelines where operational logic and execution are cleanly separated.
Inter-Agent Communication Protocols.
When multiple reasoning entities interact—whether as explicit agents or modular system components—the choice of communication protocol becomes critical. Fixed protocols (e.g., ANP (Chang et al., 2025), A2A (Project, 225), ACP (Team, 2025a)) offer stability and predictability, while learnable communication channels allow adaptive coordination but require stronger regularization to avoid emergent pathologies. Standardized communication interfaces support composability, reproducibility, and compatibility with external workflow engines.
6.6 Broader Social Impact
The growing deployment of autonomous, agentic LLM systems raises broader societal considerations that increasingly shape research priorities. This subsection highlights five cross-cutting impact areas which merit sustained attention as agentic capabilities continue to advance.
Dual-Use Risks.
The deployment of Agentic RL lowers the barrier for misuse, notably through “sleeper agent” behaviors where models appear aligned during training but activate concealed harmful policies in deployment (Hubinger et al., 2024). This deceptive alignment often persists despite SFT, RLHF, and adversarial training, as models—particularly those utilizing chain-of-thought—learn to distinguish evaluation contexts from operation. To govern such hazards across domains like Cybersecurity, CBRN, and Autonomous Replication and Adaptation (ARA), OpenAI’s Preparedness Framework establishes a four-tier risk assessment structure (OpenAI, 2025b). However, the framework faces criticism for permissive thresholds, discretionary evaluation protocols, and static assessments that fail to account for post-deployment capability evolution (Coggins et al., 2025).
Environmental Sustainability.
Large-scale RL is substantially more resource-intensive than SFT due to rollout generation, long-horizon reasoning, and iterative decision steps. Agentic systems further increase training- and deployment-time carbon footprints as interactions unfold over multi-stage workflows (Gardner et al., 2025). Sustainable practices include hardware-aware quantization and resource-efficiency–focused methods. HAQ searches for optimal layer-wise bitwidths under hardware constraints (Wang et al., 2019), and HERO derives low-bit quantization policies for efficient inference using RL-based optimization (Zhang et al., 2025u). Other recent work develops environmental evaluation benchmarks (Wu et al., 2025j).
Labor Market Implications.
The shift from token-level assistance to autonomous workflow execution positions agentic systems as increasingly strong substitutes for humans in a variety of knowledge-intensive tasks. Code agents have demonstrated the ability to perform debugging, patching, and repository-level issue resolution in SWE benchmarks (Jimenez et al., 2024; Liu et al., 2023c). GUI and web agents similarly automate interactive desktop and browser workflows as shown in OSWorld and WebArena evaluations (Xie et al., 2024; Zhou et al., 2024b). Economic analyses indicate that such end-to-end automation may disproportionately affect entry-level or routine cognitive roles, raising concerns about skill ladder erosion and labor displacement (Eloundou et al., 2023; Brynjolfsson et al., 2023). These trends highlight broader socioeconomic implications, especially for labor markets that may be increasingly exposed to automation.
Bias Amplification.
RLHF and RLAIF exacerbate societal biases and ideological sycophancy by overfitting to annotator preferences (Casper et al., 2023). Despite surface-level politeness, these models intensify covert discrimination and gender stereotypes, particularly in multi-turn agentic settings (Barnhart et al., 2025a). Furthermore, standard optimization risks collapsing minority preference modes (Xiao et al., 2025). Mitigation strategies address both reward and policy levels. Techniques include fairness-aware reward learning (Swamy et al., 2024; Ouyang et al., 2025), MaxMin-RLHF for heterogeneous groups (Chakraborty et al., 2024), and diversity-preserving objectives like DivPO to prevent mode collapse (Xiao et al., 2025; Wang et al., 2023a; Lanchantin et al., 2025). Complementary approaches involve pluralistic annotator pools and Constitutional AI (Santurkar et al., 2023; Bai et al., 2022), evaluated via benchmarks like CrowS-Pairs and dialect-sensitive tests (Barnhart et al., 2025b).
Evaluation Contamination.
Static benchmarks like HumanEval and SWE-bench suffer from data contamination, causing inflated scores and illusory robustness (Banerjee et al., 2024). In agentic settings, this encourages overfitting to environmental quirks rather than generalizable reasoning. Addressing these limitations, recent work prioritizes dynamic, contamination-resistant benchmarks, including LiveCodeBench (Jain et al., 2025b), LiveSearchBench (Zhou et al., 2025c), LiveTradeBench (Yu et al., 2025b), and LiveBench (White et al., 2025). Combined with adversarial frameworks like Breakpoint (Hariharan et al., 2025), these approaches prevent test-gaming and offer rigorous assessments of out-of-distribution performance.
Collectively, these broader-impact considerations reinforce the importance of coupling methodological advances in Agentic RL with safety, sustainability, fairness, and robustness principles. As agentic systems advance toward broader deployment, understanding and mitigating these societal effects will remain an important direction for future research.
@@ 7 Conclusion
This survey has charted the emergence of Agentic Reinforcement Learning (Agentic RL), a paradigm that elevates LLMs from passive text generators to autonomous, decision-making agents situated in complex, dynamic worlds. Our journey began by formalizing this conceptual shift, distinguishing the temporally extended and partially observable MDPs (POMDPs) that characterize Agentic RL from the single-step decision processes of conventional RL for LLMs. From this foundation, we constructed a comprehensive, twofold taxonomy to systematically map the field: one centered on core agentic capabilities (planning, tool use, memory, reasoning, self-improvement, perception, etc.) and the other on their application across a diverse array of task domains. Throughout this analysis, our central thesis has been that RL provides the critical mechanism for transforming these capabilities from static, heuristic modules into adaptive, robust agentic behavior. By consolidating the landscape of open-source environments, benchmarks, and frameworks, we have also provided a practical compendium to ground and accelerate future research in this burgeoning field.